OpenStudy (anonymous):

FIND THE FOCUS OF THE PARABOLA. -x^2-4x+3y-13=0 ************please help*********

OpenStudy (ranga):

Separate out the y and write the function as y = .... Write the equation of the parabola in vertex form. Once you do that we can find the focus.

OpenStudy (anonymous):

yes i did that... 3y=x^2+4x=13 then i divided everything by 3 and it got confusing @ranga

OpenStudy (ybarrap):

Put in this form: \(y = ax^2 + bx + c \), where, $$ a = \frac{1}{4p}; \ \ b = \frac{-h}{2p}; \ \ c = \frac{h^2}{4p} + k; \ \ $$ The vertex is at the coordinate <h,k>: $$ (x - h)^2 = 4p(y - k) \ $$ So solve the 1st set of equations for \(p,h\) and \(k\). To find the focus, add \(<0,p>\) to the vertex, \(<h,k>\)

OpenStudy (ybarrap):

See http://en.wikipedia.org/wiki/Parabola

OpenStudy (anonymous):

tamilanda...

OpenStudy (anonymous):

yes i get it but i get confused only with this problem b/c of the fraction @ybarrap @mastrblastr

OpenStudy (ranga):

-x^2-4x+3y-13=0 3y = x^2 + 4x + 13 divide by 3 y = 1/3(x^2 + 4x + 13) We did not divide each term by 3 because when completing square we want the x^2 coefficient to be 1 and so we will leave 1/3 as a factor outside. Can you complete the square of (x^2 + 4x + 13) ?